OpenStudy (anonymous):

Divide and express the result in the form P(x) = D(x)Q(X) + R(X) (4x^3 + 2x^2 + 3x + 4) divided by (x+4)

hero (hero):

Basically what you have to do is this:

hero (hero):

Start with this: 4x^2(x + 4) You'll end up with 4x^3 + 16x^2 Which means since you have 16x^2 you need to subtract 14x^2 to compensate since 16x^2 - 14x^2 = 2x. So next you'll do this: 4x^2(x + 4) - 14x^2(x + 4) since -14x(x + 4) = -14x^2 - 56x You'll need to add 59x to compensate since -56x + 59x = 3x. Then you'll have 4x^2(x + 4) - 14x^2(x + 4) + 59(x + 4) 59(x + 4) = 59x + 236 But now you need to subtract 232 to compensate since you want 236 - 232 = 4 So finally you'll have 4x^2(x + 4) - 14x^2(x + 4) + 59(x + 4) - 232 And then factor the x + 4 to get (4x^2 - 14x^2 + 59)(x + 4) - 232

hero (hero):

Good luck understanding all of that.

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

can i ask another question?

OpenStudy (anonymous):

here it is: one root of the equation x^3+x^2-2 = 0 is 1. what are the other two roots? aka how do i find these other two roots? thanks

hero (hero):

If 1 is a root of x^3 + x^2 - 2, then x = 1 is a solution and x - 1 = 0 is a factor.

OpenStudy (anonymous):

so generaly, what are the baseline rules or steps that i need to follow to find roots in this type of problem?

OpenStudy (anonymous):

im not sure if i understand what you just sent ^

hero (hero):

If you have a cubic equation and you know one of the roots, you can reduce the cubic to a quadratic equation. And all quadratic equations are factorable.

hero (hero):

Since x - 1 is a factor, re-write x^3 + x^2 - 2 as follows: x^3 - x^2 + 2x^2 - 2x + 2x - 2 = x^2(x - 1) + 2x(x - 1) + 2(x - 1) = (x^2 + 2x + 2)(x - 1)

OpenStudy (anonymous):

thanks, how did you get from x^3 + x^2 - 2 to x^3 - x^2 + 2x^2 - 2x + 2x - 2

hero (hero):

Check your messages.

hero (hero):

Well, put it this way x^2 = -x^2 + 2x^2 0 = -2x + 2x That probably still doesn't help you understand it though. But I can explain it to you in vyew.

OpenStudy (anonymous):

yes, please. it would really help if you explained it to me in vyew. thanks. ill check my messages